OpenStudy (anonymous):

Solve for y using the addition principle: = -1/8 + y= -3/4

OpenStudy (mathstudent55):

Since -1/8 is being added to y, you need to add its additive inverse, so 1/8 and -1/8 will add up to zwero, and you'll have y alone on the left side. The addition principle says you must add the same thing to both sides of an equation, so you add 1/8 to the right side too. The additive inverse of -1/8 is +1/8. Add 1/8 to both sides: 1/8 + (-1/8) + y = -3/4 + 1/8 y = -3/4 + 1/8 Find a common denominator for 4 and 8, and add the fractions on the right side/

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

You're welcome.